OpenStudy (anonymous):

|x+4|≥13 write as interval notation

OpenStudy (anonymous):

\[x + 4 \le -13\]OR\[x + 4 \ge 13\]

OpenStudy (anonymous):

Solve each inequality. \[x \le -17\]OR\[x \ge 9\]

OpenStudy (anonymous):

(-infinity, -17] U [9, infinity)